Math Across Time and Cultures

Fall Semester. Mathematics Over Time

How did mathematics grow out of human beings’ attempts to understand the natural world? Where did the mathematical topics you learned in high school come from? Are there distinctive features of specifically mathematical knowledge? In this course we will study some of the development of arithmetic, geometry, and algebra over time. We will take a close look at some of the most important sources, including the Elements of Euclid, and trace their influence on how mathematics is understood and presented today. One focus will be the development of the idea of proof and its role in mathematics.

Spring Semester. Mathematics Across Cultures

Were the Islamic artists who created the geometric mosaics of the Alhambra in Granada, Spain “doing mathematics?” To what extent are mathematical ideas a part of all human cultures? In this course, we will examine examples of ways of thinking about the world, works of art, crafts, games, and so forth from several different African, Asian, Native American, and European cultures. These ideas and objects will share many features with concepts developed in Western mathematics and we will analyze and study them using that mathematics. But did the creators of these ideas and objects think about them in the same way? And are there implications for how mathematics should be taught and learned in a multicultural world?